the value of personal computer is decreasing linearly over time.Two points indicate its price at two different times:
after one year Price=Rs.45000.
after two year Price=Rs.40.000
(1)Determine a linear equation representing this information.
(2)Predict when the computer will be worthless.
1
Expert's answer
2012-05-31T08:08:46-0400
Let's construct a line crossing points (1;45000) and (2,40000):

(t-1)/(2-1) = (P-45000)/(40000-45000);

t-1 = 9-P/5000;

P = 5000(10-t).

Here P is price and t is time (in years).

(2)Predict when the computer will be worthless.

The computer will be worthless when price will be equal to zero:

5000(10-t) = 0 ==> t = 10.

So, the computer will be worthless after 10 years.